employees of a discount appliance store receive an additional 20% off of the lowest price on an item. IF an employee purchase a dishwasher during a 15% off sales, HOW MUCH WILL PAY IF THE DISWASHER ORIGINALLY COST $450?
Answer by edjones(8007)

450-450*.15=450-67.50=382.5
382.5-382.5*.2=382.5-76.5=$306. Answer